OpenStudy (anonymous):

Prove that if a and c are odd integers, ab + bc is even for every integer b. @ganeshie8 @myininaya @UnkleRhaukus @RadEn @robtobey @chmvijay @Luigi0210 @eliassaab @genius12 @Vincent-Lyon.Fr @Preetha @Kainui @ehuman @wolfe8 @lucaz @INeedHelpPlease? @khadeeja @thadyoung @A_clan @tester97 @Andras @Confusionist @linh412986 @BlackLabel @leozap1 @JMark @Microrobot @zacharyf @Rubio101 @NaomiBell1997 @MeganEdward @theballer225 @link,zoro @Goldenmoon @CrayolaCrayon_ @Cutefriendzoned @coffeeismylife @christopher_miller420 @soumyamohan @zubhanwc3 @Ahmed47

OpenStudy (anonymous):

ab + bc = b(a+c). Let a = 2m+1 for \(\bf m \ge 0\) and let c = 2n + 1 for \(\bf n \ge 0\). Then we have: b(a+c) = b( 2m+1 + 2n + 1) = b(2m + 2n + 2) = 2[b(m + n + 1)]. Hence ab + bc is even for all integers b where a and c are odd integers.

OpenStudy (anonymous):

Something actually harder? like im not joking, im bored af..u can literally do most of these problems in ur head -.-

OpenStudy (anonymous):

Haha I"m sorry but this is the junk that I'm doing for class next semester! Good thing I'm studying ahead of time (: 4.0 here I come again! Haha, gimme a sec.

OpenStudy (anonymous):

Btw your answer lies here: http://en.wikipedia.org/wiki/Continuous_function_%28topology%29#Continuous_functions_between_topological_spaces It's a basic proof, as @eliassaab has demonstrated.

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

damn wrong thread lol